please help me solve this question A car travels at 60km per hour. How far does it travel in 5 hour 20 minutes
Found 3 solutions by Boreal, amfagge92, MathTherapy:
Answer by Boreal(15235)

5h20m is 5 1/3 hours
60*5 1/3=300+20=320 km.

Answer by amfagge92(93)

320km, speed=Distance/time taken
5h20min=5+1/3=5.3hrs
60*5.33=319.8~320km

Answer by MathTherapy(10551)


please help me solve this question A car travels at 60km per hour. How far does it travel in 5 hour 20 minutes
Never make math problems more complicated than they already are, if they are!
Correct way:
You see, you don't have to change the fraction to a decimal.
It only makes the problem more complicated, and sometimes more time-consuming, and most times, this leads to an incorrect answer.